What is the percentage water in alum [K(H2O)6Al(H2O)6(SO4)2]?


1
Expert's answer
2022-03-08T14:05:06-0500

The molecular weight of alum is 474.39

It contains 12 molecules of water or: 12 * 18.02 = 216.24

% H2O = 216.24 / 474.39 = 0.4558 or 45.58%